Solve:(1+1/2)(1+1/4)(1+1/16)(1+1/256).....infinity?

Grade:10

2 Answers

Vivek
6 years ago
Answer is infinite. These you will learn in sequence and series when you will learn divergent and convergent sequence. There you will find its answer here it is difficult to explain this will come in college 1st year
sahil
6 years ago
see the answer is not infinity its convergence it can be calculated by limit.Here is ansIf you multiply it by (1-1/2) that is by 1/2 you see (1-1/2)(1+1/2)(1+1/4)....=(1-1/4)(1+1/4)(1+1/16)...=(1-1/16)(1+1/16)(1+1/256)... you will keep going and get finally (1-1/infinity) taking limit (1-1/infinity)=1 now divide by 1/2 as you mutiplied early givind answer 2.[Hence answer is 2]